OpenStudy (anonymous):

Does this expression x2 + 24x + 144 fit into any special pattern, or not? If it does, which one(s)?

ganeshie8 (ganeshie8):

\( x2 + 24x + 144 \) \(= x^2 + 2(12)x + 12^2 \) \(= (x+12)^2 \) it can be writen as sum of two things squared ! we can use the formula \(a^2+2ab+b^2 = (a+b)^2\) to factor the expression.

ganeshie8 (ganeshie8):

@0mgl0la does it help.. ? the special pattern is the formula itself...